39 liters of water at 20 ºC are mixed with 21 liters of water at 60 ºC. What is the temperature of the mixture?
1
Expert's answer
2021-04-30T11:29:16-0400

According to thermal equilibrium equation:


"Q_\text{lost}+Q_\text{acquired}=0,\\
cm_1(t_1-t)+cm_2(t_2-t)=0,\\\space\\
t=\frac{(m_1t_1+m_2t_2)}{m_1+m_2}=34°\text C."